# Accelerated Spatio-Temporal Bayesian Modeling for Multivariate Gaussian Processes

### v1

@@ Abstract
Multivariate Gaussian processes (GPs) offer a powerful probabilistic framework to represent complex interdependent phenomena. They pose, however, significant computational challenges in high-dimensional settings, which frequently arise in spatial-temporal applications. We present DALIA, a highly scalable framework for performing Bayesian inference tasks on spatio-temporal multivariate GPs, based on the methodology of integrated nested Laplace approximations. Our approach relies on a sparse inverse covariance matrix formulation of the GP, puts forward a GPU-accelerated block-dense approach, and introduces a hierarchical, triple-layer, distributed memory parallel scheme. We showcase weak scaling performance surpassing the state-of-the-art by two orders of magnitude on a model whose parameter space is 8$\times$ larger and measure strong scaling speedups of three orders of magnitude when running on 496 GH200 superchips on the Alps supercomputer. Applying DALIA to air pollution data from northern Italy over 48 days, we showcase refined spatial resolutions over the aggregated pollutant measurements.
@@
Spatio-Temporal Modeling, Large-scale Bayesian Inference, Distributed Memory Computin
@@ II-A Univariate Gaussian Processes
For $\bm{x}=[x_{1},x_{2},\dots,x_{n}]$, each variable $x_{j}$ is associated with a spatial or spatio-temporal location $s_{j}$, such that $x_{j}=x(s_{j})$. The sparse symmetric positive definite precision matrix $\bm{Q}_{p}(\bm{\theta})$ is parametrized by the hyperparameters $\bm{\theta}$ that govern the spatial or spatio-temporal dependencies. For brevity, we will implicitly assume this dependency on $\bm{\theta}$ from now on and omit in the notation. Being able to evaluate Eq. 1 efficiently for fixed values of $\bm{\theta}$ and $\bm{x}$ is essential to the overall inference process and arises repeatedly for different parameter configurations. It requires the factorization of $\bm{Q}_{p}$ to compute its determinant, denoted by ||$\cdot$|| ⋅ |, from the diagonal entries of the Cholesky factor. The size of $\bm{x}$ and $\bm{Q}_{p}$ grows with increasing numbers of time steps or refined spatial resolutions.
@@ IV DALIA Framework Innovation
$N$ Total size of the BTA matrix, equal to $nb+a$.
@@ IV-B Multivariate Model Formulation
where $\bm{Q}^{i}$ refers to the precision matrix (fixed effects included) of each univariate process as defined in Eq. 2. Both the multivariate prior $\bm{Q}^{nv}_{p}$ and the conditional precision matrix $\bm{Q}_{c}^{nv}$ are constructed using Eq. 11, when respectively relying on $\bm{Q}^{i}_{p}$ and $\bm{Q}^{i}_{c}$. While the joint precision matrices $\bm{Q}^{n_{v}}_{p/c}$ remain sparse, however, and as shown in Fig. 2b), the overall BT/BTA structure is not maintained. The dimensions of the joint matrices are the sum of the dimensions of the univariate precision matrices, i.e. dim$(\bm{Q}^{n_{v}})=N=n_{v}(n_{s}n_{t}+n_{r})$, where $n_{v}$ is the number of univariate processes, $n_{s}$ the spatial mesh size, $n_{t}$ the number of time steps and $n_{r}$ the number of fixed effects for each univariate process. For simplicity, we assume them to be the same. If their dimensions differed one would additionally require projection mappings between the different meshes for their joint matrix assembly. Each of the $\bm{Q}^{i}_{p}$ consist of the sum of sparse Kronecker products arising from the spatial and temporal mesh discretization [25] and has a separate set of unknown hyperparameters which was omitted in the notation above for readability.
@@ V-A Evaluation Methodology and Hardware
In Table IV we present an overview of the dimensions of the different models and datasets used for our performance analyses. We conducted our GPU experiments at the Swiss national computing center (CSCS) on the Alps supercomputer equipped with NVIDIA GH200 superchips (Grace CPU and Hopper GPU), using the NVIDIA Collective Communications Library (NCCL) at the solver-level (mostly intranode communication) and CUDA-AWARE-MPI otherwise. We run the R-INLA package on the Intel Xeon Platinum 8470 “Sapphire Rapids” CPU (2TB partition) of the Fritz supercomputer at the Erlangen National High Performance Computing Center (NHR@FAU). We used the sparse direct solver PARDISO backend for R-INLA and empirically determined the most favorable nested OpenMP thread configuration for R-INLA, following the guidelines provided in [43]. We chose $\mathcal{S}_{1}=9$, $PARDISO_{\text{omp}}=8$ for case study $MB_{1}$ and $\mathcal{S}_{1}=8$, $PARDISO_{\text{omp}}=8$ for case study $WA_{1,2}$ and $SA_{1}$. Here $\mathcal{S}_{1}$ represents the number of groups performing the finite difference function evaluation in parallel, and $PARDISO_{\text{omp}}$ the number of threads per sparse solver instance within each group $\mathcal{S}_{1}$.
@@ V-C Solver Microbenchmarks
Figure 6: Weak scaling results of the DALIA framework on the Grace-Hopper GPUs of the CSCS-Alps supercomputer, references runtimes for the R-INLA package achieved on the Fritz supercomputer of the FAU computing center. $a)$ Weak scaling through temporal discretization of a trivariate coregional model, from $2$ time-steps ($1$ GPU) to $512$ time-steps ($248$ GPUs). $b)$ Weak scaling through spatial mesh refinement, from $72$ mesh nodes ($1$ GPU) to $4485$ nodes ($496$ GPUs). $c)$ Spatial meshes at different refinement levels on northern region of Italy.
@@ VI Application to Air Pollution Dataset
We apply our framework to a multivariate spatio-temporal coregionalization model on air pollution, jointly modeling PM2.5, PM10, and O3[52, 12]. We used hourly pollutant measurements provided by the Copernicus Atmosphere Monitoring Service (CAMS) [53] with a spatial resolution of $0.1^{\circ}$ ($\approx 10\text{km}$). We aggregated the data to daily resolution for analysis, covering a $122{,}350\,\text{km}^{2}$ region in northern Italy with 4210 spatial locations and 48 days of observations, starting from January 1st, 2022, see also $AP_{1}$ in Table IV. Using our framework we perform spatial downscaling to a resolution of $0.02^{\circ}$ ($\approx 2\text{km}$), increasing spatial detail 25-fold. The results of the downscaling process of the multivariate GP for ozone (O3) concentration are presented in Fig. 8. The first row shows the coarse satellite measurements, while the second row displays the disaggregated estimated mean. The first column represents the averaged data over the entire time period and reveals consistently high average ozone levels in oceanic areas. A time-resolved analysis reveals that the averaged data over the entire time period fails to represent the behavior of individual days, including unusually low concentrations in the eastern region on January 1 (day 1) and high concentrations in the western region on February 1 (day 32). This behavior is, however, accurately captured by our spatio-temporal approach, underpinning the need for efficient implementations that allow for spatio-temporal instead of only spatial multivariate modeling. Our approach also provides posterior estimates for the models’ interpretable parameters, such as the effect of elevation on pollutant concentrations. Specifically, a 1 km increase in elevation is associated with reductions of $0.45~{}\mu g/m^{3}$ in PM2.5, with the 0.025 and 0.975 quantiles estimated at $-0.51$ and $-0.40$, respectively; a reduction of $0.55~{}\mu g/m^{3}$ in PM10, with corresponding quantiles at $-0.62$ and $-0.49$; and an increase of $1.27~{}\mu g/m^{3}$ in O3, with quantiles of $1.20$ and $1.34$. In addition, we obtain the following linear correlations between pollutants, revealing a strong positive correlation ($0.97$) between PM2.5 and PM10, and moderate negative correlations with O3: $-0.61$ for PM2.5 and $-0.63$ for PM10.
@@ Acknowledgments
This work was supported by the Swiss National Science Foundation (SNSF) under grant $\mathrm{n^{\circ}}$ 209358 (QuaTrEx), and by the Platform for Advanced Scientific Computing in Switzerland (BoostQT). We acknowledge the scientific support and HPC resources from CSCS under projects sm96 and lp16 as well as the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project 80227.

### v2

@@ Abstract
[…]
@@
Spatio-Temporal Modeling, Large-scale Bayesian Inference, Distributed Memory Computing
@@ II-A Univariate Gaussian Processes
For $\bm{x}=[x_{1},x_{2},\dots,x_{n}]$, each variable $x_{j}$ is associated with a spatial or spatio-temporal location $s_{j}$, such that $x_{j}=x(s_{j})$. The sparse symmetric positive definite precision matrix $\bm{Q}_{p}(\bm{\theta})$ is parametrized by the hyperparameters $\bm{\theta}$ that govern the spatial or spatio-temporal dependencies. For brevity, we will implicitly assume this dependency on $\bm{\theta}$ from now on and omit in the notation. Being able to evaluate Eq. 1 efficiently for fixed values of $\bm{\theta}$ and $\bm{x}$ is essential to the overall inference process and arises repeatedly for different parameter configurations. It requires the factorization of $\bm{Q}_{p}$ to compute its determinant, denoted by $|\cdot|$, from the diagonal entries of the Cholesky factor. The size of $\bm{x}$ and $\bm{Q}_{p}$ grows with increasing numbers of time steps or refined spatial resolutions.
@@ IV DALIA Framework Innovation
NN Total size of the BTA matrix, equal to $nb+a$.
@@ IV-B Multivariate Model Formulation
[…]
@@ V-A Evaluation Methodology and Hardware
In Table IV we present an overview of the dimensions of the different models and datasets used for our performance analyses. We conducted our GPU experiments at the Swiss national computing center (CSCS) on the Alps supercomputer equipped with NVIDIA GH200 superchips (Grace CPU and Hopper GPU), using the NVIDIA Collective Communications Library (NCCL) at the solver-level (mostly intranode communication) and MPI otherwise. We run the R-INLA package on the Intel Xeon Platinum 8470 “Sapphire Rapids” CPU (2TB partition) of the Fritz supercomputer at the Erlangen National High Performance Computing Center (NHR@FAU). We used the sparse direct solver PARDISO backend for R-INLA and empirically determined the most favorable nested OpenMP thread configuration for R-INLA, following the guidelines provided in [43]. We chose $\mathcal{S}_{1}=9$, $PARDISO_{\text{omp}}=8$ for case study $MB_{1}$ and $\mathcal{S}_{1}=8$, $PARDISO_{\text{omp}}=8$ for case study $WA_{1,2}$ and $SA_{1}$. Here $\mathcal{S}_{1}$ represents the number of groups performing the finite difference function evaluation in parallel, and $PARDISO_{\text{omp}}$ the number of threads per sparse solver instance within each group $\mathcal{S}_{1}$.
@@ V-C Solver Microbenchmarks
Figure 6: Weak scaling results of the DALIA framework on the Grace-Hopper GPUs of the CSCS-Alps supercomputer, references runtimes for the R-INLA package achieved on the Fritz supercomputer of the FAU computing center. $a)$ Weak scaling through temporal discretization of a trivariate coregional model, from 22 time-steps ($1$ GPU) to $512$ time-steps ($248$ GPUs). $b)$ Weak scaling through spatial mesh refinement, from $72$ mesh nodes (11 GPU) to $4485$ nodes ($496$ GPUs). $c)$ Spatial meshes at different refinement levels on northern region of Italy.
@@ VI Application to Air Pollution Dataset
We apply our framework to a multivariate spatio-temporal coregionalization model on air pollution, jointly modeling PM2.5, PM10, and O3[52, 12]. We used hourly pollutant measurements provided by the Copernicus Atmosphere Monitoring Service (CAMS) [53] with a spatial resolution of $0.1^{\circ}$ ($\approx 10\text{km}$). We aggregated the data to daily resolution for analysis, covering a $122\,350\,\text{km}^{2}$ region in northern Italy with 4210 spatial locations and 48 days of observations, starting from January 1st, 2022, see also $AP_{1}$ in Table IV. Using our framework we perform spatial downscaling to a resolution of $0.02^{\circ}$ ($\approx 2\text{km}$), increasing spatial detail 25-fold. The results of the downscaling process of the multivariate GP for ozone (O3) concentration are presented in Fig. 8. The first row shows the coarse satellite measurements, while the second row displays the disaggregated estimated mean. The first column represents the averaged data over the entire time period and reveals consistently high average ozone levels in oceanic areas. A time-resolved analysis reveals that the averaged data over the entire time period fails to represent the behavior of individual days, including unusually low concentrations in the eastern region on January 1 (day 1) and high concentrations in the western region on February 1 (day 32). This behavior is, however, accurately captured by our spatio-temporal approach, underpinning the need for efficient implementations that allow for spatio-temporal instead of only spatial multivariate modeling. Our approach also provides posterior estimates for the models’ interpretable parameters, such as the effect of elevation on pollutant concentrations. Specifically, a 1 km increase in elevation is associated with reductions of $0.45~\mu g/m^{3}$ in PM2.5, with the 0.025 and 0.975 quantiles estimated at 0.51-0.51 and 0.40-0.40, respectively; a reduction of $0.55~\mu g/m^{3}$ in PM10, with corresponding quantiles at 0.62-0.62 and 0.49-0.49; and an increase of $1.27~\mu g/m^{3}$ in O3, with quantiles of $1.20$ and $1.34$. In addition, we obtain the following linear correlations between pollutants, revealing a strong positive correlation ($0.97$) between PM2.5 and PM10, and moderate negative correlations with O3: 0.61-0.61 for PM2.5 and 0.63-0.63 for PM10.
@@ Acknowledgments
This work was supported by the Swiss National Science Foundation (SNSF) under grant $\mathrm{n^{\circ}}$ 209358 (QuaTrEx), grant $\mathrm{n^{\circ}}$ 200021 (NumESC), and by the Platform for Advanced Scientific Computing in Switzerland (BoostQT). We acknowledge the scientific support and HPC resources from CSCS under projects sm96 and lp16 as well as the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project 80227.